# Advances and Open Challenges in Federated Learning with Foundation Models

### v1

@@ II Terminology
This section provides an overview of FMs and FL, outlining the fundamental concepts and historical development trajectories that have propelled their current prominence in the field.
@@ II-B Federated Learning
The typical FL process involves several key steps. Initially, a global model is distributed to all participating FL clients from the FL server. Each client then trains the ML model on its local data to derive an updated local model. These model updates are subsequently sent back to the FL server, where they are aggregated to produce an updated global model. This cycle is repeated until convergences or specific performance criteria are met. A widely used framework in FL is Federated Averaging (FedAvg) [25], renowned for its ability to aggregate model updates with minimal communication overhead. This is particularly crucial in FL settings, where a potentially large number of participants may have limited communication bandwidth [26].
To achieve meaningful real-world impact, FL must address several intricate challenges, including non-IID (independently and identically distributed) data across clients, system heterogeneity, and scalability [13]. Non-IID data can introduce biases into models by favoring the data distribution of specific participants [27]. System heterogeneity, characterized by variations in computation and communication capabilities among client participants, can lead to uneven contributions to the model training process. Additionally, scalability concerns arise as the number of participants increases, necessitating the development of efficient algorithms and infrastructure to manage the aggregation of updates and distribution of the global model. The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
@@ III The Proposed FedFM Taxonomy
In the rapidly evolving domain of FL, the integration of FMs into federated settings, termed FedFM, represents a significant leap forward. This integration aims to leverage the capabilities of FMs and the collaborative training process of FL to enable FedFM to access privately owned decentralized data. The development of FedFM requires rethinking of several key aspects of the current FL paradigm: 1) achieving efficiency in federated training and aggregation, 2) ensuring trustworthiness, and 3) building effective incentive mechanisms. Each plays a crucial role in the successful implementation and operation of FedFM. Based on the above considerations, we propose a multi-tiered FedFM taxonomy, as illustrated in Fig. 1.
@@ III-A Efficient Federated Training and Aggregation: The Foundation of FedFM
  • Computationally Efficient FedFM: Research in this area focuses on adapting pre-trained FMs to specific tasks with minimal adjustments to the model parameters. Techniques like Parameter-Efficient Fine-Tuning (PEFT), Prompt Tuning (PT), and Instruction Tuning (IT) are possible ways to enhance FedFM computational efficiency. These methods allow for significant reductions in the computational and storage demands by fine-tuning only a small subset of the model parameters.
@@ IV Towards Efficient FedFM
11:  $\mathcal{B}\leftarrow$ (split $\mathcal{D}_{c}$ into batches of size $B$)
@@ IV-A FedFM Training and Aggregation
To grasp the challenges involved in integrating FMs into the FL process, a revision of the general FL procedure is necessary. FL typically operates under the coordination of a central server with the primary purpose of collaboratively training a Deep Learning (DL) model $\bm{\theta}$ across $N$ clients over multiple rounds, as outlined in Algorithm 1. The integration of FMs, with their considerable sizes, introduces challenges at various stages of the FL process. These challenges can be distilled into three critical research questions (highlighted in blue in Algorithm 1):
@@
    How to aggregate FMs to enhance performance? Given that one of the primary goals of FL is to improve global model performance, the challenge lies in designing aggregation approaches that are suited for the scale and complexity of FMs. Despite the large body of existing FL model aggregation techniques, few are tailored for FMs.
@@
  • Model Compression: which aims to reduce the overall size of the model being transmitted.
@@ IV-B Scaling Efforts for FedFM
We list the methods reviewed alongside with the largest base models involved in their evaluations in Table III and Table IV. The substantial difference in model sizes between existing FL research and centralized FMs might cast doubts on the feasibility and practicality of the FL techniques for FedFM, calling for further experimental studies involving large-scale FMs. Such an effort depends on the availability of frameworks that can streamline implementation. Thus, in Section VII, we analyze the support for FMs in existing FL frameworks.
@@ IV-C1 Weighted Average
As shown in Algorithm 1, at communication round $t$ the local training process in traditional FL yields a set of local models $\{\bm{\theta}_{1}^{t},\bm{\theta}_{2}^{t},\ldots,\bm{\theta}_{m}^{t}\}$. These models are transmitted to the server where they undergo an aggregation process to produce global model $\bm{\theta^{t+1}}\leftarrow\text{Aggregate}(\{\bm{\theta}_{1}^{t},\bm{\theta}_{2}^{t},\ldots,\bm{\theta}_{m}^{t}\})$. The aggregation step is pivotal in FL, as the overall performance of the FL model depends on it [52]. Consequently, the development of aggregation methods that balance efficiency with efficacy remains a central theme in FL research. The predominant approach involves computing a weighted average of the parameters from local models (i.e., the weighted average method). Prominent examples of this approach include FedAvg/FedSGD [25], FedProx [28], and FedMA [53].
To the best of our knowledge, there is currently no FL aggregation technique specifically tailored for FedFM. The prevailing practice has gravitated towards adopting the basic vanilla versions of FedAvg or FedSGD, with the latter being equivalent to FedAvg when the number of local epochs ($E$) is set to one. Table V provides a summary of the aggregation methods adopted by existing FedFM works.
@@ IV-D1 Parameter-Efficient Fine-Tuning
Under FL settings, Sun et al. [42] introduced FedPEFT, an innovative approach for fine-tuning FMs. This method freezes the majority of the model weights, focusing on adjusting a minimal subset of parameters tailored to specific downstream tasks. FedPEFT evaluates the efficiency of four distinct PEFT strategies including, Head-tuning, Bias, Adapter and Prompt-in the context of the Vision Transformers (ViT-B) with 85 million parameters. Their findings demonstrate that these targeted fine-tuning methods achieve comparable results to centralized model fine-tuning on non-IID data, while remarkably reducing communication overhead by over 99%. FedCLIP [12] also experimented with an Adaptor-based PEFT method using FedAvg-based adapter aggregation technique. Their method surpasses the performance of standard FedAvg and FedProx. FedPETuning [43] conducted an extensive benchmark analysis of adapter-based fine-tuning techniques under FL settings. Their findings indicate that both approaches, integrating additional adapters and utilizing LoRA, can achieve comparable accuracy. However, LoRA-based fine-tuning incurs only one third of the communication overhead required by methods involving extra adapters.
@@
FwdLLM [47] is designed to adapt FMs for use on mobile devices with stricter resource constraints. To address this issue, it integrates Backpropagation (BP)-free training with parameter-efficient training approaches. Rather than relying on traditional back-propagation to calculate precise gradients, the BP-free method introduces minor, self-generated perturbations to the model weights. It then evaluates how these perturbations affect model predictions compared to the original unperturbed model. If a perturbation results in predictions that are closer to the actual labels, it is considered to be guiding the model towards the global optimum. Remarkably, FwdLLM demonstrates substantial performance enhancements, achieving convergence three orders of magnitude faster and reducing the memory footprint by close to 15 times.
@@ IV-D3 Instruction Tuning
Instruction tuning is an effective method for improving the functionality and manageability of LLMs [67]. It involves additional training of FMs using pairs of instructions and the corresponding outputs, where an instruction specifies a task for a model and the output represents the expected result in accordance with the given instruction. The appeal of employing instruction tuning for FedFM stems not only from its computational efficiency, which facilitates the quick adaptation of FMs to particular domains without the need for extensive retraining or modifications to the model structure. In addition, it also effectively narrows the divergence between FMs’ inherent next-word prediction goals and the actual user intent, aligning with the aim of personalized FL [27]. This approach also leads to more consistent and predictable model behaviors. Through instruction-based fine-tuning, models are more closely guided in output generation, ensuring alignment with user expectations and enhancing overall controllability.
@@ IV-E Communication Efficient FedFM
The complexity of FMs directly translates to inflated communication overhead between FL clients and the FL server in FedFM, hindering efficient collaboration. Although as discussed in Section IV-D, it is not always necessary to transmit an entire FM, fine-tuning billion-parameter FMs with adapter-based methods still requires the transmission of millions of parameters. This highlights the need for communication efficient FedFM research. In this section, we highlight two primary strategies for enhancing efficiency: 1) model pruning and 2) model compression.
@@ IV-E1 Model Pruning
Model pruning aims to identify and retain only the parts of a model that are essential for a specific task. An early model pruning approach is HeteroFL [31], which aims to accommodate heterogeneous model architectures within FL by adaptively distributing sub-networks suited to the capacities of individual clients. It selects and aggregates subsets of the global model, pruning of a large model into a variety of smaller more manageable models in effect. FjORD [33] implements ordered dropout, a technique that organizes knowledge within a deep neural network in a structured and hierarchical manner. This allows for the extraction of compact sub-models without retraining. Ordered dropout enhances computational efficiency by dropping model components in sequence rather than at random, aligning with the optimization capabilities of contemporary linear algebra libraries. The approach is complemented by a self-distillation process to refine the model further.
Building on these foundations, PruneFL [35] proposed a two-stage model pruning process specifically designed for FL environments. Initially, a “warm-up” phase involves selecting a single capable and trusted client to prune the model using its local data, thereby starting the FL process with a streamlined model. In the subsequent “adaptive pruning” phase, the server periodically adjusts the model by removing or reintroducing parameters over multiple iterations. FedPM [36] adopts a strategy inspired by the lottery ticket hypothesis for model pruning. Rather than utilizing a pruned model as an initial point, it initializes a random binary mask guided by a common seed for each client. At the end of each round of FL training, clients return their binary masks to the server. The server then constructs a global model by computing a weighted average of these masks. FedTiny [37] follows a similar approach to PruneFL, but with a novel initialization step. It utilizes batch normalization values from clients’ data as the basis for selecting a shared initialization, thereby enhancing adaptability to diverse client data distributions.
@@ IV-E2 Model Compression
Deep models often operate on full precision (32-bit). This high degree of detail is not always necessary. FedPAQ [30] leverages this fact to reduce communication overhead. Note that dynamic quantization only benefits communication. Depending on the infrastructure, the model updates might have to be cast back to full precision, thereby incurring additional computational overhead. SoteriaFL [38] strikes a balance among privacy, convergence accuracy and communication efficiency. For model compression, it utilizes a straightforward algorithm, CDP-SGD, which effectively integrates communication compression with DP-SGD for enhanced efficiency.
@@ V-A2 Byzantine-Robust Local Model Aggregation
Geometrical Outlier Detection: These schemes discard local model updates which are regarded as geometric outliers, and only aggregates the remaining ones to form the global model. Some approaches [86] calculate the Euclidian distances among local model updates to determine the divergences among them, and remove those with the highest degree of heterogeneity from the majority. Others [87] analyze the parameter-wise geometrical divergence, removing the largest and smallest values of the same parameter of all model updates, and only aggregating the remainders to produce the final corresponding parameter in the global model.
@@
Other Hybrid Schemes: Research works have emerged [91, 92, 93] in an attempt to form hybrids schemes to take advantages of the aforementioned by combining them together. They usually update a benchmark model on a central clean dataset, and lower the weights assigned to local model updates that significantly deviate from the benchmark (or even discard them completely) during aggregation. Although these methods have achieved superior robustness, they have also inherited the limitations from the aforementioned categories of approaches which make them incompatible with the FedFM settings.
@@ VI Towards Incentive Mechanisms for FedFM
The remarkable success of FL depends on clients (a.k.a., data owners) actively engaging in the model training process. In reality, data owners might hesitate to join FL without proper compensation, especially if it involves significant commitment of local resources for training FMs. Thus, developing robust incentive mechanisms is imperative to motivating clients to participate in the FedFM training process, while deterring misbehaviours through punitive measures. Table VI provides an overview of existing incentive mechanisms in FL.
@@ VI-D FedFM Contribution Evaluation
Contribution evaluation in FedFM is more challenging compared to the current approaches for centralized FMs or traditional FL. Firstly, different from centralized FM evaluations, FedFM cannot directly assess the quality of the training data used due to privacy protection requirements. Secondly, FedFM evaluation is limited by the resource constraints in terms of local computation and communication power of the FL clients involved. Thirdly, compared to traditional FL, FedFM evaluation can involve a large number of sequences, resulting in insufficient test coverage [162].
There are a number of existing evaluation methods for traditional FL and centralized FMs, but few for FedFM. In this section, we discuss existing evaluation methods and envision promising research directions for designing new FedFM evaluation protocols. We present approaches which cover relevant works to evaluate the model performance, explain the prediction results, and provide insights into the contributions made by individual participants or data samples.
@@ VI-D1 Contribution Evaluation for Traditional FL
The performance of the global FL model largely depends on the quality of the local data. Client and sample contribution evaluation helps the FL server understand this by tracing back to local data, and is important to model aggregation. FL client and sample contribution evaluation methods can be divided into two main categories: 1) Shapley value (SV)-based evaluation and 2) influence-based evaluation (Table VII).
@@
Shapley Value-based Evaluation: SV-based evaluation methods provide insights into FL clients’ data through evaluating their contributions to the performance of the final FL model. Existing SV-based evaluation methods mostly focus on improving computational efficiency, while maintaining SV estimation accuracy. The original SV method is prohibitively expensive since the required utility function evaluation grows exponentially with the number of FL participants. Existing FL evaluations based on efficient SV calculation can be divided into two categories: 1) accelerating within-round evaluations, and 2) reducing the number of rounds of sub-model evaluation. For the first approach, instead of re-training them from scratch, gradient-based SV [163] and local embedding-based SV [164] methods have been proposed to reconstruct sub-models.
@@
For the second approach, three main methods have been proposed. The first evaluates every possible sub-model as in the original SV setting based on gradient-based estimation [130, 165]. The second method leverages randomly sampled permutation evaluation [166, 167] which produces the estimated SV of an FL participant as its expected contribution. However, since the number of selected permutations is fixed, potentially important permutations might be overlooked, resulting in inaccurate estimation. Thus, the third approach [168, 169] combines the guided Monte Carlo sampling technique and the within-round and between-round truncation techniques to prioritize sub-model permutations based on their importance.
@@
Another branch studies the problem of secure SV calculation under malicious settings. In [170], a two-server secure SV calculation protocol was proposed, which leverages a hybrid scheme to avoid ciphertext-ciphertext multiplications. Another work [171] proposed a group-based SV computation scheme leveraging a blockchain-based secure framework with secure aggregation to protect participants’ data privacy.
@@
Influence-based Evaluation: Although SV-based FL client and sample contribution evaluation consider the complex dependencies among clients, they are generally highly expensive to compute. Influence-based FL contribution evaluation methods have been proposed to efficiently evaluate the contribution of FL clients and their local data samples on FL model performance. Existing influence-based methods can be divided into two main branches. The first [172, 173] perturbs clients or their local samples to retrain FL models, and uses the difference in performance (e.g., test loss, test accuracy) between the new model and the original one to approximate client or sample contribution. However, since this method requires retraining of FL models on all clients’ datasets or individual samples, the evaluation procedure is prohibitively expensive [174, 175].
@@
The second category uses influence function methods [175] that leverage the second-order optimization technique to avoid the expensive retraining. An early approach leveraging influence functions in client contribution evaluation is Fed-influence [176]. To measure the influence of a client, Fed-influence sums up all sample influence values since there is an additive property of the influence function when measuring changes in test predictions [177]. However, this approach requires clients to calculate the inverse of the Hessian matrix and transmit it, which incurs large computation and communication overhead. To reduce the overhead of influence approximation, emerging methods [174, 178, 179, 180] leverage the Hessian vector product (HVP) to approximate the influence values. They are capable of achieving the linear operation costs, making them promising for practical adoption.
@@
Model-Agnostic Evaluation: These methods treat an FL model as a black-box and attempts to measure the relevance of each feature to the learning task to filter out irrelevant ones. They leverage various statistical measures (e.g., mutual information, Gini-impurity, F-statistics [181, 182, 183]) to compute per-feature relevance scores. Another branch of methods use efficient gradient-based SV estimation approaches for feature contribution evaluation [172, 184]. However, they methods require the FL server to know all the IDs of clients’ features, which might violate clients privacy and make them unsuitable for practical vertical FL applications.
Model Specific Evaluation: The utilization of attention-based evaluation method enables the server to interpret which specific parts of inputs are leveraged by the global FL model. In [185], a hierarchical attention mechanism is proposed which develops task-specific attentions to access personal feature correlations. Besides, a temporal attention layer is designed to evaluate cross-client temporal correlations at the FL server level. The final visualization of the attention weights can be used to determine which features the global model focuses for individual predictions. Another work, Flames2Graph [186], offers a personalized evaluation approach for the multivariate time series FL classification problem. It extracts and visualizes the essential sequences that highly activate network neurons to capture the temporal dependencies among them. Furthermore, [187, 188] proposed a federated feature selection approach, which introduces a Gaussian stochastic dual-gate based on the $l_{0}$ constraints to efficiently and privately approximate the probability of a feature being selected.

### v2

@@ II Terminology
This section aims to provide an overview of FMs and FL, outlining the fundamental concepts and historical development trajectories that have propelled their current prominence in the field.
@@ II-B Federated Learning
The typical FL process involves several key steps. Initially, a global model is distributed to all participating FL clients from the FL server. Each client then trains the ML model on its local data to derive an updated local model. These model updates are subsequently sent back to the FL server, and then are aggregated to update the global model. Such cycle is repeated until convergences or specific performance criteria are met. A widely used framework in FL is Federated Averaging (FedAvg) [25], renowned for its ability to aggregate model updates with minimal communication overhead. This is particularly crucial in FL settings, where a potentially large number of participants may have limited communication bandwidth [26].
To achieve meaningful real-world impact, FL must address several intricate challenges, including non-independently and identically distributed (non-IID) data across clients, system heterogeneity, and scalability [13]. Non-IID data can introduce biases into models by favoring the data distribution of specific participants [27]. System heterogeneity, characterized by variations in computation and communication capabilities among client participants, can result in uneven contributions to the model training process. Additionally, scalability concerns arise as the number of participants increases, necessitating the development of efficient algorithms and infrastructure to manage the aggregation of updates and distribution of the global model. The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
@@ III The Proposed FedFM Taxonomy
In the rapidly evolving domain of FL, the integration of FMs into federated settings, termed FedFM, represents a significant leap forward. This integration aims to leverage the capabilities of FMs and the collaborative training process of FL to enable FedFM to access privately owned decentralized data. The development of FedFM requires rethinking of several key aspects of the current FL paradigm: 1) achieving efficiency in federated training and aggregation, 2) ensuring trustworthiness, and 3) building effective incentive mechanisms. Each part plays a crucial role in the successful implementation and operation of FedFM. Based on the above considerations, we propose a novel and multi-tiered FedFM taxonomy, as illustrated in Fig. 1.
@@ III-A Efficient Federated Training and Aggregation: The Foundation of FedFM
  • Computationally Efficient FedFM: Research in this area focuses on adapting pre-trained FMs to specific tasks with minimal adjustments to the model parameters. Techniques like Parameter-Efficient Fine-Tuning (PEFT), Prompt Tuning (PT), and Instruction Tuning (IT) are possible ways to enhance FedFM computational efficiency. These methods allow for significant reductions in the computational and storage demands by only fine-tuning a small subset of the model parameters.
@@ IV Towards Efficient FedFM
11:  $\mathcal{B}\leftarrow$ (split $\mathcal{D}_{c}$ into batch of size $B$)
@@ IV-A FedFM Training and Aggregation
To grasp the challenges involved in integrating FMs into the FL process, a revision of the general FL procedure is necessary. FL typically operates under the control and decision-making of the central server with the primary purpose of collaboratively training a Deep Learning (DL) model $\bm{\theta}$ across $N$ clients over multiple rounds, as outlined in Algorithm 1. The integration of FMs, with their considerable sizes, introduces challenges at various stages of the FL process. These challenges can be distilled into three critical research questions (highlighted in blue in Algorithm 1):
@@
    How to aggregate FMs to enhance performance? Given that one of the primary goals of FL is to improve the global model performance, the challenge lies in designing aggregation approaches that are suited for the scale and complexity of FMs. Despite the large body of existing FL model aggregation techniques, few are tailored for FMs.
@@
  • Model Compression: which aims to decrease the total number of the model parameters being transmitted.
@@ IV-B Scaling Efforts for FedFM
We list the methods reviewed alongside with the largest base models involved in their evaluations in Table III and Table IV. The substantial difference for model sizes between existing FL research and centralized FMs might cast doubts on the feasibility and practicality of the FL techniques for FedFM, calling for further experimental studies involving large-scale FMs. Such an effort depends on the availability of frameworks that can streamline implementation. Thus, in Section VII, we analyze the support for FMs in existing FL frameworks.
@@ IV-C1 Weighted Average
As shown in Algorithm 1, the local training process of traditional FL yields a set of local models $\{\bm{\theta}_{1}^{t},\bm{\theta}_{2}^{t},\ldots,\bm{\theta}_{m}^{t}\}$ at communication round $t$. These models are transmitted to the server where they undergo an aggregation process to produce global model $\bm{\theta^{t+1}}\leftarrow\text{Aggregate}(\{\bm{\theta}_{1}^{t},\bm{\theta}_{2}^{t},\ldots,\bm{\theta}_{m}^{t}\})$. The aggregation step is pivotal in FL, as the overall FL model performance relies on it [52]. Consequently, the development of aggregation methods that balance efficiency with efficacy remains a central theme in FL research. The predominant approach involves computing a weighted average of the parameters from local models (i.e., the weighted average method). Prominent examples of this approach include FedAvg/FedSGD [25], FedProx [28], and FedMA [53].
To the best of our knowledge, currently, there is not FL aggregation technique specifically tailored for FedFM. The prevailing practice has gravitated towards adopting the basic vanilla versions of FedAvg or FedSGD, with the latter being equivalent to FedAvg when the number of local epochs ($E$) is set to one. Table V provides a summary of the aggregation methods adopted by existing FedFM works.
@@ IV-D1 Parameter-Efficient Fine-Tuning
Under FL settings, Sun et al. [42] introduced FedPEFT, an innovative approach for fine-tuning FMs. This method freezes the majority of the model weights, focusing on adjusting a minimal subset of parameters tailored to specific downstream tasks. FedPEFT evaluates the efficiency of four distinct PEFT strategies including, Head-tuning, Bias, Adapter and Prompt-in the context of the Vision Transformers (ViT-B) with 85 million parameters. Their findings demonstrate that these targeted fine-tuning methods achieve comparable results to centralized model fine-tuning with non-IID data, while remarkably reducing communication overhead by over 99%. FedCLIP [12] also experimented with an Adaptor-based PEFT method using adapter aggregation even based on FedAvg, showing the better performance than standard FedProx and FedAvg. FedPETuning [43] conducted an extensive benchmark analysis of adapter-based fine-tuning techniques under FL settings. Their findings indicate that both approaches, integrating additional adapters and utilizing LoRA, can achieve comparable accuracy. However, LoRA-based fine-tuning incurs only one third of the communication overhead required by methods involving extra adapters.
@@
FwdLLM [47] is designed to adapt FMs for use on mobile devices with stricter resource constraints. To address this issue, it integrates Backpropagation (BP)-free training with parameter-efficient training approaches. Rather than relying on traditional back-propagation to calculate precise gradients, the BP-free method introduces minor and self-generating perturbations to these ML model parameters. It then evaluates how these perturbations affect model predictions compared to the original unperturbed model. If a perturbation results in predictions that are closer to the actual labels, it is considered to be guiding the model towards the global optimum. Remarkably, FwdLLM demonstrates substantial performance enhancements, achieving convergence three orders of magnitude faster and reducing the memory footprint by close to 15 times.
@@ IV-D3 Instruction Tuning
Instruction tuning is an effective method for improving the functionality and manageability of LLMs [67]. It involves additional training of FMs using pairs of instructions and the corresponding outputs, where an instruction specifies a task for a model and the output represents the expected result in accordance with the given instruction. The appeal of employing instruction tuning for FedFM stems not only from its computational efficiency, which facilitates the quick adaptation of FMs to particular domains. In this way, there is no need for extensive re-training or modifications to the model structure. In addition, it also effectively narrows the divergence between FMs’ inherent next-word prediction goals and the actual user intent, aligning with the aim of personalized FL [27]. This approach also leads to more consistent and predictable model behaviors. Through instruction-based fine-tuning, models are more closely guided in output generation, ensuring alignment with user expectations and enhancing overall controllability.
@@ IV-E Communication Efficient FedFM
The complexity of FMs directly translates to inflated communication overhead between FL clients and the FL server in FedFM, hindering efficient collaboration. Although as discussed in Section IV-D, it is not always necessary to transmit an entire FM, fine-tuning billion-parameter FMs by adapter still requires the transmission of millions of parameters. This highlights the need for communication efficient FedFM research. In this section, we highlight two primary strategies for enhancing efficiency: 1) model pruning and 2) model compression.
@@ IV-E1 Model Pruning
Model pruning aims to identify and retain only the essential parts of the model for a specific purpose. An early model pruning approach is HeteroFL [31], which aims to accommodate heterogeneous model architectures within FL by adaptively distributing sub-networks suited to the capacities of individual clients. It selects and aggregates subsets of the global model, pruning of a large model into a variety of smaller more manageable models in effect. FjORD [33] implements ordered dropout, a technique that organizes knowledge within a deep neural network in a structured and hierarchical manner. This allows for the extraction of compact sub-models without retraining. Ordered dropout enhances computational efficiency by dropping model components in sequence rather than at random, aligning with the optimization capabilities of contemporary linear algebra libraries. The approach is complemented by a self-distillation process to refine the model further.
Building on these foundations, PruneFL [35] proposed a two-stage model pruning process specifically designed for FL environments. Initially, a “warm-up” phase involves selecting a single capable and trusted client to prune the model using its local data, thereby starting the FL process with a streamlined model. In the subsequent “adaptive pruning” phase, the server periodically adjusts the model by removing or reintroducing parameters over multiple iterations. FedPM [36] adopts a strategy inspired by the lottery ticket hypothesis for model pruning. Rather than utilizing a pruned model as an initial point, it initializes the random binary mask guided by the common seed for all clients. At the end of each round of FL training, clients return their binary masks to the server. The server then constructs a global model by computing a weighted average of these masks. FedTiny [37] follows a similar approach to PruneFL, but with a novel initialization step. It utilizes batch normalization values from clients’ data as the basis for selecting a shared initialization, thereby enhancing adaptability to diverse client data distributions.
@@ IV-E2 Model Compression
Deep models often operate with full precision (32-bit), but in practice such high degree for computation may be not always necessary. FedPAQ [30] leverages this fact to reduce communication overhead. Note that dynamic quantization only benefits communication. Depending on the infrastructure, the model updates might need to cast back to full precision, thereby incurring additional computational burden. SoteriaFL [38] strikes a balance among privacy, convergence accuracy and communication efficiency. For model compression, it utilizes a straightforward algorithm, CDP-SGD, which effectively integrates communication compression with DP-SGD for enhanced efficiency.
@@ V-A2 Byzantine-Robust Local Model Aggregation
Geometrical Outlier Detection: These schemes discard local model updates which are regarded as geometric outliers, and only aggregates the remaining ones to form the global model. Some approaches [86] calculate the Euclidian distances among local model updates to determine the divergences among them, and remove those with the highest degree of heterogeneity from the majority. Others [87] analyze the parameter-wise geometrical divergence, removing the smallest and largest values of the same parameter of all model updates, and only aggregating the remainders to produce the final corresponding parameter in the global model.
@@
Other Hybrid Schemes: Research works have emerged [91, 92, 93] in an attempt to form hybrids schemes to take advantages of the aforementioned by combining them together. These works often update a benchmark model with a central clean dataset, and lower the weights assigned to local model updates that significantly deviate from the benchmark (or even discard them completely) during aggregation. Although these methods have achieved superior robustness, they have also inherited the limitations from the aforementioned categories of approaches which make them incompatible with the FedFM settings.
@@ VI Towards Incentive Mechanisms for FedFM
The remarkable success of FL depends on clients (a.k.a., data owners) actively engaging during the training process. In reality, data owners might hesitate to join FL without proper compensation, especially if it involves significant commitment of local resources for training FMs. Thus, developing robust incentive mechanisms is imperative to motivating clients to participate in the FedFM training process, while deterring misbehaviours through punitive measures. Table VI provides an overview of existing incentive mechanisms in FL.
@@ VI-D FedFM Contribution Evaluation
Contribution evaluation in FedFM is more challenging compared to the current approaches for centralized FMs or traditional FL. Firstly, different from centralized FM evaluations, FedFM cannot directly assess the quality of the training data used due to privacy protection requirements. Secondly, FedFM evaluation is limited by the resource constraints in terms of communication power and local computation of the FL clients involved. Thirdly, compared to traditional FL, FedFM evaluation can involve a large number of sequences, resulting in insufficient test coverage [162].
There are a number of existing evaluation methods for traditional FL and centralized FMs, but few for FedFM. In this section, we discuss existing evaluation methods and envision promising research directions for designing new FedFM evaluation protocols. We present approaches which cover relevant works to evaluate the model performance, analysis the decision results, and provide insights into the contributions of individual participants and datasets.
@@ VI-D1 Contribution Evaluation for Traditional FL
The performance of the global FL model highly relies on the quality of the local dataset. Client and sample contribution evaluation can help the FL server analysis this through tracing back to local data, and is important to model aggregation. FL client and sample contribution evaluation methods can be divided into two main categories: 1) Shapley value (SV)-based evaluation and 2) influence-based evaluation (Table VII).
@@
Shapley Value-based Evaluation: SV-based evaluation methods provide insights into FL clients’ data via evaluating their contributions to the performance of the final FL model. Existing SV-based evaluation methods mostly focus on improving computational efficiency, while maintaining accuracy performance of the estimated SV. The original SV method is prohibitively expensive since the required utility function evaluation grows exponentially with the number of FL participants.
@@
Existing FL evaluations based on efficient SV calculation can be divided into two categories: 1) accelerating within-round evaluations, and 2) decreasing the number of rounds of sub-model evaluation. For the accelerating within-round evaluation approach, instead of re-training them from scratch, gradient-based SV [163] and local embedding-based SV [164] methods have been proposed to reconstruct sub-models. For the decreasing the number of rounds of sub-model evaluation approach, three popular methods have been proposed. The first evaluates every possible sub-model within the original SV setting based on gradient-based estimation [130, 165]. The second method leverages randomly sampled permutation evaluation [166, 167] which produces the estimated SV of an FL participant as its expected contribution. However, since the number of selected permutations is fixed, potentially important permutations may be overlooked, leading to inaccurate estimation. Thus, the third approach [168, 169] combines the within-round and between-round truncation approach and the guided Monte Carlo sampling and in order to prioritize sub-model permutations based on their importance.
@@
Another branch studies the problem of secure SV calculation under malicious settings. In [170], a two-server secure SV calculation protocol was designed, which leverages a hybrid scheme to avoid ciphertext-ciphertext multiplications. Another work [171] designed a group-based SV computation scheme leveraging a blockchain-based secure aggregation framework in order to protect participants’ data privacy.
@@
Influence-based Evaluation: Although SV-based FL client and sample contribution evaluation consider the complex dependencies among clients, they are generally highly expensive to compute. Influence-based FL contribution evaluation methods have been proposed to efficiently evaluate the contribution of FL clients and their local data samples on FL model performance. Existing influence-based methods can be divided into two main branches. The first [172, 173] perturbs clients or their local samples to retrain FL models, and uses the difference in performance (e.g., test loss, test accuracy) between the new and the original model to approximate client or sample contribution. However, since this method requires retraining of FL models on all clients’ datasets or individual samples, the evaluation procedure is prohibitively expensive [174, 175].
@@
The second category uses influence function methods [175] that leverage the second-order optimization technique to avoid the expensive retraining. An early approach leveraging influence functions in client contribution evaluation is Fed-influence [176]. To measure the influence of a client, Fed-influence sums up all sample influence values since there is an additive property of the influence function when measuring changes in test predictions [177]. However, this approach requires clients to calculate the inverse of the Hessian matrix and transmit it, which incurs large computation and communication overhead. To reduce the overhead of influence approximation, emerging methods [174, 178, 179, 180] leverage the Hessian vector product to approximate the influence values. They are capable of achieving the linear operation costs, making them promising for practical adoption.
@@
Model-Agnostic Evaluation: These methods consider an FL model as a black-box and attempts to measure the relevance of each feature to the learning task to filter out irrelevant ones. They leverage various statistical measures (e.g., mutual information, F-statistics, Gini-impurity[181, 182, 183]) to compute per-feature relevance scores. Another branch of methods use efficient gradient-based SV estimation approaches for feature contribution evaluation [172, 184]. However, they methods require the FL server to familiarize all the IDs of clients’ features, which might violate clients privacy and make them unsuitable for practical vertical FL applications.
Model Specific Evaluation: The utilization of attention-based evaluation method enables the server to interpret which specific parts of inputs are leveraged by the global FL model. In [185], a hierarchical attention mechanism is proposed which develops task-specific attentions to access personal feature correlations. Besides, a temporal attention layer is designed to evaluate cross-client temporal correlations at the FL server level. The final visualization of the attention weights is used to determine which features the global model focuses for individual predictions. Another work, Flames2Graph [186], offers a personalized evaluation approach for the multivariate time series FL classification issue, which can extract and visualize the essential sequences that highly activate network neurons to capture the temporal dependencies among them. Furthermore, [187, 188] proposed a federated feature selection method, which introduces a Gaussian stochastic dual-gate based on the $l_{0}$ constraints to efficiently and privately approximate the probability of selected feature.